• Log in or sign up for an account here: https://www.sythe.org/login

easy 10m for stakers

easy 10m for stakers
It's one of the most commonly use betting 'strategies' you will see around the duel arena.

Very easy to lose out on big cash, unless you know how to run your stakes.
 
easy 10m for stakers
your correct this has already been tried many many times, but it doesnt always work like you say, because you could run into someone who has a great ammount of PID, or could just as easily run into a scammer and get scammed for one of those ammounts, lowering the chances of you losing all 7 in a row, in my opinion and with my experiences, i'd say just stay away from staking all together, even though it is the best way to make cash, it's also the best way to lose cash.
 
easy 10m for stakers
Best off just staking what your going to risk, gambling methods like this have been around for 100's of years and have proven to not work.
 
easy 10m for stakers
Except if you lose all of that in a row, you're screwed. Trust me it happens, I've lost 12 stakes in a row before, whilst doing the 2x method.
 
easy 10m for stakers
There is a 50% chance you will lose all 7 stakes. They are each independent factors.

I'm not sure if you know about this but lol...

0.5x0.5x0.5x0.5x0.5x0.5x0.5 = 0.0078125 Or, in otherwords the probability you will lose all 7 stakes
 
easy 10m for stakers
Jack, the stakes are each separate incidents. You losing one has no effect on the next one whatsoever. Each time you start a stake, it's a 50% chance you'll win, and a 50% chance you'll lose. Just as one can lose 10 stakes in a row someone can also win 10 stakes in a row.
 
easy 10m for stakers
I do that all the time :p I start with 1m stakes xd
 
easy 10m for stakers
This is like 100m /hr, even more if you got around ~50b

???
profit
 
easy 10m for stakers
Bad idea. I've lost 3B from this. Got cleaned twice.

The odds of losing are much highest than that. And you cannot determine your odds. Because if you put yourself up with someone that has lost 10 stakes in a row and did near 0s in his last few fights, your odds aren't 50/50, the odds are on their side. I know someone that lost 15 stakes in a row.

Don't bother with this is my only tip for you. If you feel lucky, go ahead and stake a few mil. But don't look at this as a way to make billions, because the odds aren't always in your favor even when you think it may be.
 
easy 10m for stakers
This is one of the most generic method of staking, but some people also like taking risks every now and then. My general method would be to do that until I've 5-10xed my cash, then go about half of what I've made in.. Even if I did lose, I still had 2x of my original amount.

I'm also one of those unlucky people who lost 13 times in a row doing this and ended up losing over 1.5B due to my luck.. I believe it was like a .13% chance of something like that happening ;3
 
easy 10m for stakers
It's a simple method that was used buy gamblers but is now outdated. Runescape really comes down to lucky though you win some you lose so, I guess you could go big when your down a few stakes and small when your on a hot streak but ... Runescape is Runescape and sometimes you will just get Jagex'd
Wrong, there is no such thing as hot streak.

There is a 50% chance you will lose all 7 stakes. They are each independent factors.
Wrong.

I'm also one of those unlucky people who lost 13 times in a row doing this and ended up losing over 1.5B due to my luck.. I believe it was like a .13% chance of something like that happening ;3
Losing 13 times in a row would be 0,0122% chance.



Can somebody tell on average how long does it take to find an opponent and fight(maxed boxing)?
 
easy 10m for stakers
easy 10m for stakers
Mathematically, you have nearly 0 percent chance of losing 7 in a row, but there is a slim chance and in the long run (I imagine many stakers like to stake a lot) it will happen. And when it does this method will make them lose their bank. There is NO betting method that will statistically work. For anyone reading this I would try finding out methods such as flicking atk and def, 1st hit, to get the advantage, or finding people with lower stats, because betting methods to overcome 50 50 odds are a farce.
 
easy 10m for stakers
you have the concepts wrong its like in genetics

when a homozygous breeds with a homozygous recessive
50% of the offspring should be heterozygous and 50% should be homozygous

that doesn't mean if the first offspring is heterozygous the second will by homozygous recessive

as previous offspring do not affect the 50% ratio for the next offspring

same with staking just because you lose a fight doesn't mean you will win the next due to the 50% chance of winning each individual fight
 
easy 10m for stakers
i think this is what most stakers do anyways :p
 
easy 10m for stakers
This doesn't work at all. I've lost over 8x in a row countless times staking. Call it bad luck but it's happened to me frequently.
 
Runescape Gamblingrunescape gambling
Back
Top